OpenStudy (anonymous):

Kiana built a scale model of a school that is under construction. On her model, there is 5.6 millimeters between two of the windows on the front of the school. The scale that Kiana used to make the model was 2.8 millimeters = 3 feet. What will the actual distance between the windows be after the school is built? A.0.2 feet B. 6 feet C. 5 feet D. 16.8 feet

OpenStudy (anonymous):

Converting between mm and feet is not funny but luckily enough we got nice numbers. 2.8 mm = 3 feet 2 * 2.8 mm = 2 * 3 feet 5.6 mm = 6 feet